# Yaroslav Aksenov, Ian Maksimov, Nikita Balagansky, Daniil Gavrilov Tinkoff

### v2

@@ 4.1 Experimental Setup
Tasks
We evaluate each training configuration on two datasets: Anthropic-HH [^1] https://huggingface.co/datasets/Anthropic/hh-rlhf (Bai et al.,, 2022) and Reddit TL;DR summarization [^2] https://huggingface.co/datasets/UCL-DARK/openai-tldr-summarisation-preferences (Stiennon et al.,, 2020). The Anthropic-HH dataset contains 160,800 training and 8,552 validation examples of pair preferences for SFT and preference learning. The Reddit TL;DR dataset includes 73,396 training and 26,198 validation examples, post-duplicate removal, of human preference judgments on model-generated summaries. We select only the uniquely preferred summaries for SFT, resulting in 41,947 training and 11,941 validation examples.
Models
We employ a range of Pythia models (Biderman et al.,, 2023), with sizes of 2.8B, 6.9B, and 12B, serving as pre-trained base models. An SFT policy checkpoint trains similarly to Rafailov et al., (2023); Liu et al., 2024b on preferred texts for each dataset. We explore two main update strategies, each adaptable to different base alignment methods (e.g., DPO, IPO, KTO) as outlined in Section 2. (1) Soft Update: This strategy applies a weighting factor $\alpha$, with experiments conducted for $\alpha\in[0.0,1.0]$, to progressively merge the current policy with its reference policy at each training step. Notably, the TR variants (TR-DPO, TR-IPO, and TR-KTO) become equivalent to their base methods (DPO, IPO, and KTO respectively) when $\alpha=0$. The notation for soft update is based on the base method, such as TR-DPOα, TR-IPOα, and TR-KTOα. (2) Hard Update: This strategy involves distinct experiments where the reference model updates at fixed intervals $\tau$ to evaluate the efficacy of varying update frequencies. The notation for the hard update method also depends on the base method and is similarly denoted as TR-DPOτ, TR-IPOτ, and TR-KTOτ. Further details on the experimental setup and hyperparameters are given in Appendix A.
Evaluation
@@
Following established approaches (Rafailov et al.,, 2023; Liu et al., 2024b, ), we employ a comprehensive evaluation framework to assess the performance of various TR methods’ configurations against the corresponding original baselines.
AutoSxS evaluation: We employ the AutoSxS framework, using ‘GPT-4-0125-preview‘ as a proxy for human evaluators (detailed prompt information can be found in Appendix H), to analyze preferences across 500 samples from the test set. This comparison includes various configurations of TR-DPO, TR-IPO, TR-KTO, and their respective traditional counterparts, DPO, IPO, and KTO, maintaining consistent generation parameters: temperature set to 1.0, top_k at 40, top_p at 0.7, and max_new_tokens at 512.
Human-centric metrics evaluation: To estimate the Probability of Improvement (PoI) (Agarwal et al.,, 2021), inspired by our proposed update strategies, we employ Human-Centric (HC) metrics for direct and absolute response assessment. These metrics, which include coherence, correctness, level of detail, helpfulness, and harmlessness, are scored on a scale from 1 to 10 and are analogous to those evaluated by Hu et al., (2023). The choice of HC metrics over Reward Models (RM) is driven by the significant KL divergence from the standard fine-tuning (SFT) policies of TR methods, potentially rendering new samples Out-Of-Distribution (OOD) for RMs and leading to inaccurate assessments. A specialized subset of 150 test samples, specifically selected for their provocative content within the Helpful and Harmless dataset, was examined. In this assessment, ‘GPT-4-0125-preview‘ serves as a proxy for human judgment, with guiding prompts provided in Appendix H.3.
Model dynamics and policy divergence: To investigate the influence of the proposed methods on KL divergence and generation diversity, we measure Self-BLEU (Zhu et al.,, 2018) and the KL divergence between the original and learned policies using the full Anthropic-HH test set for the Pythia 2.8B model. Additionally, we compare these metrics with the previously described mean HC metrics. This analysis allows for a detailed exploration of how our proposed methods modify the dynamics compared to simply changing the $\beta$, and evaluates the trade-off between alignment and diversity (Kirk et al.,, 2023; Ahmadian et al.,, 2024; Wang et al.,, 2023).
(a) TR-DPO vs DPO on the Anthropic-HH dataset.
(b) TR-DPO vs. DPO on the Reddit TL;DR dataset.
@@ 4.2 Performance Comparison on the Two Tasks
For a comprehensive exploration of update strategies across the entire range of $\alpha$ and $\tau$, we employed the TR-DPO approach using the Pythia 2.8B model. The base model was configured with $\beta=0.05$, demonstrating an optimal trade-off between Human-Centric metrics and KL divergence, as detailed in Section 4.4. This setup allows for a systematic assessment of the impact of each strategy under varied conditions, effectively comparing the effects of different weighting factors $\alpha$ and update intervals $\tau$.
Figures 3(a) and 3(b) illustrate that both the soft and hard update strategies of TR-DPO enhance performance when compared to the traditional DPO method for the Pythia 2.8B model on the Anthropic-HH and Reddit TL;DR datasets. TR-DPO with $\alpha$ values between 0.5 and 0.6 consistently outperforms settings with other $\alpha$ values. Conversely, the benefits of TR-DPOτ become more pronounced as $\tau$ increases from 64 to 512.
For both datasets, the parameters $\alpha=0.6$ and $\tau=512$ for soft and hard updates, respectively, pass the Fisher statistical test with the Pythia 2.8B model size. Detailed results for the GPT comparison are presented in Appendix Table 3.
We assert that these parameters are near-optimal for most scenarios, and can be applied to various methods, tasks, and model sizes. This makes it possible to keep the computation budget the same for the proposed methods and baselines, maintaining the purity of the experiments in our work and facilitating hyperparameter tuning for further practical use of our methods. The results of the experiments are presented in Table 1 and include three methods (TR-DPO, TR-IPO, and TR-KTO), two tasks (Anthropic HH and Reddit TLDR), and three model sizes (2.8B, 6.9B, and 12B). For each method, an optimal $\beta$ was selected and used in training the models (see Section 4.4 for more details). Examples of the generations of the models can be found in Appendix G.
@@
4.3 Detailed Human-Centric Analysis
Figure 4 presents standardized absolute scores for HC metrics, including coherence, correctness, level of detail, helpfulness, and harmlessness, as evaluated on the Anthropic-HH dataset. For TR-DPO, the optimal $\alpha$ values are between 0.5 and 0.6. Performance degradation is observed as $\alpha$ deviates from this optimal range, particularly at higher values such as 0.8, where the model exhibits training instability and a tendency to generate repetitive words. A similar trend is observed for frequent updates when $\tau$ is equal to 64 (see Appendix Figure 7).
Building on these findings, the Probability of Improvement (PoI) metric (Agarwal et al.,, 2021) further validates these effects. Employing bootstrap sampling to generate confidence intervals and medians, PoI illustrates the superiority of our TR-DPO method over the traditional DPO baseline. Significant improvements are documented in Figures 5(a) and 5(b), particularly for $\alpha$ values of 0.5 and 0.6 for soft updates, and $\tau$ intervals of 512 and 1024 for hard updates. These results, based on the Pythia 2.8B model, demonstrate that the confidence intervals do not overlap the 0.5 probability threshold, thereby underscoring the statistical significance of the TR method enhancements.
@@ 4.3 Detailed Human-Centric Analysis
While the results show a noticeable improvement in metrics such as coherence, correctness, helpfulness, and harmlessness, there is no improvement in level of detail. This can be explained by the fact that this criterion depends on the length of generated texts. As previously noted in Park et al., (2024), DPO tends to increase the text length. In Appendix Figure 16, we show that TR-DPO is less subject to this problem, as it tends to produce shorter texts for the comparable levels of KL divergence.
@@
Similarly, both TR-IPO and TR-KTO show statistically confirmed enhancements over their corresponding baselines. A detailed analysis of these methods, including specific $\alpha$ and $\tau$ parameters, is provided in Appendix C.
4.4 Why Just Lowering $\beta$ Is Not Enough
It is clear that updating the reference policy of the trained one increases the KL divergence between the trained policy and the original SFT. Equation 1 shows that the same effect can be achieved by reducing the $\beta$ coefficient of the constraint on the KL divergence. The same parameter is also present in equations 2, 3 and 4. Figure 6(a) demonstrates the dependency of the average value of the HC metrics and the KL divergence with the SFT on the $\beta$ parameter for the DPO, TR-DPOα and TR-DPOτ methods, using the Pythia 2.8B model on the Anthropic-HH dataset. We note that the best value of HC metrics is achieved at $\beta=0.05$, which we use for DPO in the experiments from Sections 4.2 and 4.3. The value of this parameter for IPO and KTO was chosen the same way (see the graphs in Appendix D). One notable result is that, after reaching a peak value, HC metrics for DPO fall faster than for its modified versions, allowing the TR methods to achieve better results.
@@ 4.4 Why Just Lowering $\beta$ Is Not Enough
As highlighted by Wang et al., (2023), models with higher alignment tend to produce less diverse generations. The dependency of HC metrics on Self-BLEU (Zhu et al.,, 2018) is shown in the Appendix Figure 15, and is similar to the dependency on KL divergence. Based on these graphs, we can affirm that the TR methods show higher values of HC metrics at the same level of response diversity (or KL divergence).
Figure 6: (a) The relationship between KL divergence and HC mean value for DPO and TR-DPO ($\alpha=0.6$, $\tau=512$) across different $\beta$ values. The best HC value for DPO occurs at $\beta=0.05$, but goes into decline after. TR-DPO methods extend further in KL divergence without a rapid decline in HC. (b) Gradient scale variations for DPO, TR-DPO with a soft update ($\alpha=0.6$), and TR-DPO with a hard update ($\tau=512$) at $\beta=0.05$ using the Pythia 2.8B model. The soft update achieves a gradual increase, while the hard update acts as a gradient reset. See Section 4.4 for details.
To understand this behavior, consider the gradient dynamics of DPO and the update strategies of both TR-DPO versions during the training process. The gradient of the loss function, specified in Equation 2, is as follows:
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{\text{ref}})=\\ &=-\beta\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\sigma\left(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\right)\nabla_{\theta}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right].\end{split}$ (8)
The absolute values of term $\sigma\left(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\right)$ act as a weighting factor, signifying the disparity between the outcomes of the trained policy and the reference model. For instance, the gradient magnitude becomes small when the probability assigned to the preferred outcome $y_{w}$ by the trained policy is higher than that of the reference model and conversely for the less preferred outcome $y_{l}$.
The gradient scales are depicted in Figure 6, smoothed for clarity. We hypothesize that there is a relationship between these observed gradient scales and the training dynamics. Specifically, a higher $\alpha$ in soft updates and a lower $\tau$ in hard updates both result in increased gradient scales, which may be indicative of greater training instability (see Appendix D.3 for more details). This has the potential to cause significant deviations from the stability of the SFT policy, leading to degradation. Conversely, updating the reference policy allows for more deviation from the SFT policy while maintaining stable performance compared to the standard DPO method. Similar behavior is observed for TR-IPO and TR-KTO as presented in Appendix D, Figures 12(b) and 13(b).
@@
Thus, simply lowering the $\beta$ coefficient is not sufficient to reach the optimal results. By updating the reference policy, TR methods achieve higher metric values and maintain stable performance compared to standard alignment techniques, indicating the importance of incorporating such update strategies to enhance training dynamics.

### v3

@@ 4.1 Experimental Setup
4.1.1 Tasks
We evaluate our training configurations on both task-specific datasets, following prior works (Liu et al., 2024b; Zhu et al., 2024; Liu et al., 2024a), and on general benchmarks, similar to Tunstall et al. (2023); Meng et al. (2024).
Task-Specific Datasets: For specialized evaluations, the Anthropic-HH [^1] https://huggingface.co/datasets/Anthropic/hh-rlhf (Bai et al., 2022) dataset, which focuses on dialogue alignment where preferred responses are selected based on their helpfulness and harmlessness. For the summarization task, we employ the Reddit TL;DR summarization [^2] https://huggingface.co/datasets/UCL-DARK/openai-tldr-summarisation-preferences (Stiennon et al., 2020) dataset, training models to generate concise and accurate summaries of long social posts.
General Benchmarks: For broader, general-purpose evaluations, we use the UltraChat-200k [^3] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k (Ding et al., 2023) dataset, designed to train models’ ability to follow instructions in open-domain conversations. Additionally, the UltraFeedback [^4] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized (Cui et al., 2023) dataset provides a binarized preference framework, useful for aligning models across various domains in an offline setting, making it suitable for training and evaluating general-purpose assistants.
A summary of the dataset sizes is provided in the Appendix D.
4.1.2 Models
Pythia Models: For task-specific experiments, we use Pythia (Biderman et al., 2023) models with 2.8B, 6.9B, and 12B parameters. Following Rafailov et al. (2023); Liu et al. (2024b), we obtain an SFT policy checkpoint by training on the preferred texts for each dataset.
Llama3 Models: For general benchmarks, we employ Llama3 (AI@Meta, 2024) models with 8B parameters in two distinct settings, as described by Tunstall et al. (2023); Meng et al. (2024): Base Setting: Llama3-Base (8B) is trained on the UltraChat-200k dataset to obtain the SFT policy, which is then aligned using the specified methods on the UltraFeedback preference dataset. Instruct Setting: Llama3-Instruct (8B) generates 8 responses per prompt from the UltraFeedback dataset. The best and worst responses are selected using PairRM [^5] https://huggingface.co/llm-blender/PairRM (Jiang et al., 2023) to form adaptation preference pairs for alignment methods.
4.1.3 Update Strategies
[…]
Further details on the experimental setup and hyperparameters are given in Appendix D.
4.1.4 Evaluation
@@
We employ a comprehensive evaluation framework, following established approaches (Rafailov et al., 2023; Lee et al., 2024; Tunstall et al., 2023; Meng et al., 2024), to assess the performance of various TR method configurations against their original baselines.
(a) Anthropic-HH
(b) Reddit TL;DR
@@
AutoSxS Evaluation: Utilizing the AutoSxS framework (Liu et al., 2024b; Rafailov et al., 2023) with ‘GPT-4-0125-preview‘ as a proxy for human evaluators (detailed prompts in Appendix L), we analyze preferences across 500 samples from the test set. This includes various configurations of TR-DPO, TR-IPO, TR-KTO, and their traditional counterparts.
General Benchmark Evaluation: We evaluate our models on AlpacaEval 2 (Li et al., 2023) and Arena-Hard (Li et al., 2024), following the evaluation setups in Tunstall et al. (2023) and Meng et al. (2024). In both settings, ‘GPT-4-1106-preview‘ is used as the judge model.
Policy Divergence and Overoptimization Analysis: To investigate the impact of our proposed methods on KL divergence and overoptimization, we measure the KL divergence between the original SFT policy and the learned policies using the full Anthropic-HH test set for the Pythia 2.8B model. We also compute Human-Centric (HC) metrics—such as coherence, correctness, level of detail, helpfulness, and harmlessness—scored on a scale from 1 to 10, similar to those evaluated by Hu et al. (2023), and calculate the Probability of Improvement (PoI) (Agarwal et al., 2021). HC metrics were computed for a specialized subset of 150 test samples with provocative content from the Helpful and Harmless dataset; we used ‘GPT-4-0125-preview‘ as a proxy judge (guiding prompts provided in Appendix L.3). We analyze the mean HC metrics versus KL divergence for classical methods compared to the proposed TR modifications at different $\beta$ values. Additionally, we compute Self-BLEU scores (Zhu et al., 2018) to evaluate how our methods affect the trade-off between alignment and diversity.
@@ 4.2 Performance Comparison on the Two Tasks
To comprehensively explore update strategies across the entire range of $\alpha$ and $\tau$, we employed the TR-DPO approach using the Pythia 2.8B model utilizing task specific datasets. The base model was configured with $\beta=0.05$, demonstrating an optimal trade-off between HC metrics and KL divergence, as detailed in Section 4.4. This setup allows for a systematic assessment of the impact of each strategy under varied conditions, effectively comparing the effects of different weighting factors $\alpha$ and update intervals $\tau$.
Figures 4(a) and 4(b) illustrate that both the soft and hard update strategies of TR-DPO enhance performance when compared to the traditional DPO method for the Pythia 2.8B model on the Anthropic-HH and Reddit TL;DR datasets. TR-DPO with $\alpha$ values between 0.5 and 0.6 consistently outperforms settings with other $\alpha$ values. Conversely, the benefits of TR-DPOτ become more pronounced as $\tau$ increases from 64 to 512. For a deeper understanding of how the behavior of the algorithms depends on $\alpha$ and $\tau$, we have plotted the scale of the loss function gradient value for several hyperparameter values. Please refer to Section H for more details.
@@
For both datasets, the parameters $\alpha=0.6$ and $\tau=512$ for soft and hard updates, respectively, pass the Fisher statistical test with the Pythia 2.8B model size. Detailed statistical test results are presented in Appendix Table 8.
We recognize that the optimal hyperparameters $\alpha$ and $\tau$ can vary across different tasks and model sizes. In our experiments, $\alpha=0.6$ and $\tau=512$ mostly demonstrated strong performance on the Pythia models for the Anthropic-HH and Reddit TL;DR datasets, as presented in Table 1 for different models sizes (2.8B, 6.9B, and 12B) and proposed methods (TR-DPO, TR-IPO, and TR-KTO). Although our ablation study (see Appendix E) shows that other hyperparameters can yield better results for specific setups, the values $\alpha=0.6$ and $\tau=512$ were used to ensure a fair comparison across methods while maintaining a consistent computational budget. Furthermore, we notice improvements in quality across the most evaluated scenarios with these selected hyperparameters. For each method, an optimal $\beta$ was selected and used in training the models (see Section 4.4 for more details). Examples of model generations can be found in Appendix K.
4.3 General Benchmarks Evaluation
Method Llama3-Base (8B) Llama3-Instruct (8B)
AlpacaEval 2 Arena-Hard AlpacaEval 2 Arena-Hard
LC(std) WR(std) WR(CI) LC(std) WR(std) WR(CI)
SFT 8.5 (0.5) 4.8 (0.6) 3.2 (-0.7, 0.8) 24.9 (0.8) 25.3 (1.3) 19.1 (-1.8, 1.8)
DPO 15.9 (0.8) 11.2 (1.0) 10.5 (-1.3, 1.6) 33.9 (0.8) 37.6 (1.4) 31.7 (-2.4, 2.4)
TR-DPOα 27.3 (0.8) 23.9 (1.2) 21.3 (-1.8, 2.2) 43.5 (0.8) 46.8 (1.5) 34.7 (-3.0, 1.6)
TR-DPOτ 27.7 (0.8) 25.7 (1.3) 20.2 (-1.8, 2.2) 42.8 (0.8) 47.2 (1.4) 32.4 (-2.3, 1.9)
IPO 15.9 (0.7) 11.5 (1.0) 8.9 (-1.1, 1.4) 36.1 (0.8) 39.2 (1.4) 31.5 (-2.0, 2.6)
TR-IPOα 29.5 (0.8) 25.4 (1.3) 19.4 (-1.8, 2.0) 43.6 (0.8) 46.9 (1.5) 33.8 (-2.6, 2.3)
TR-IPOτ 28.1 (0.8) 25.4 (1.3) 21.1 (-1.8, 2.5) 42.6 (0.8) 46.8 (1.5) 34.5 (-2.4, 2.1)
KTO 15.8 (0.8) 13.0 (1.0) 9.7 (-1.3, 1.4) 32.1 (0.9) 33.5 (1.4) 26.1 (-1.9, 2.3)
TR-KTOα 14.7 (0.7) 12.5 (1.0) 14.9 (-1.6, 1.8) 37.9 (0.9) 40.0 (1.4) 29.2 (-1.9, 2.2)
TR-KTOτ 16.5 (0.7) 13.8 (1.0) 14.3 (-1.6, 1.6) 41.3 (0.8) 41.2 (1.4) 30.9 (-2.1, 2.1)
Table 2: Results for AlpacaEval 2 and Arena-Hard experiments using $\alpha=0.8$ and $\tau=32$ for TR methods. LC and WR denote length-controlled (Dubois et al., 2024) and raw win rate, respectively. We train SFT models for Base settings on the UltraChat dataset. For Instruct settings, we use off-the-shelf models as the SFT model, as mentioned in Section 4.1. Bold values represent the best performance for each benchmark, while underlined values represent the second-best performance. See Section 4.3 for more details.
The results in Table 2 demonstrate that TR modifications consistently and statistically significantly outperform their respective base methods across both the AlpacaEval 2 and Arena-Hard benchmarks. Notably, all TR methods with hard update configurations show substantial improvements over their base counterparts.
However, for KTO, the results are mixed. While TR-KTOτ outperforms its base counterpart, the soft update version, TR-KTOα, shows slightly lower results compared to the base KTO on Llama3-Base in terms of LC win rate, though it demonstrates solid improvements on Arena-Hard. Further hyperparameter tuning, as discussed in Appendix E, yield better results for KTO configurations. For this table, the parameters $\alpha=0.8$ and $\tau=32$ were used for consistency.
4.4 Divergence and Overoptimization Analysis
Figure 5: The relationship between KL divergence and HC mean value for (a) DPO/TR-DPO, (b) IPO/TR-IPO, and (c) KTO/TR-KTO ($\alpha=0.6$, $\tau=512$) across different $\beta$ values. While for low KL values both vanilla and TR methods show similar HC values, as the KL increases, the vanilla methods start to suffer from overoptimization. In contrast, TR methods show better quality at large KL values, supporting the hypothesis from Section 3.1. See Section 4.4 for details.
As discussed in Section 3.1, overoptimization is associated with an increasing KL divergence between the trained policy $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$. In offline alignment methods, overoptimization can lead the model to assign higher probabilities to OOD data, thereby degrading performance. By setting $\beta$ to a lower value, we can regularize the DPO Hessian (Eq. 7) and prevent vanishing curvature, as $\sigma(s)$ remains closer to 0.5. Motivated by the goal of reducing overoptimization through Trust Region (TR) methods, we evaluated DPO/TR-DPO, KTO/TR-KTO, and IPO/TR-IPO on the Anthropic-HH dataset using the Pythia 2.8B model across various $\beta$ values to explore overoptimization in these methods.
@@ 4.4 Divergence and Overoptimization Analysis
To compare HC/KL dynamics, we used $\alpha=0.6$ and $\tau=512$, as PoI (Agarwal et al., 2021) (Figures 6(a) and 6(b)) shows that the confidence intervals do not overlap the 0.5 probability threshold, indicating the statistical significance [^6] Based on the PoI method as detailed in Agarwal et al. (2021). of TR method enhancements. These hyperparameters are thus appropriate for comparing the performance within this setup. Similarly, both TR-IPO and TR-KTO show statistically significant improvements over their corresponding baselines. A detailed analysis of these methods, including specific $\alpha$ and $\tau$ parameters, is provided in Appendix G.
As shown in Figure 5, TR variants of all methods are less affected by overoptimization. At equivalent levels of KL divergence, TR variants produce a higher mean HC metric, demonstrating a better Pareto front. Vanilla methods start to degrade in quality earlier, while TR methods can progress further, achieving a higher peak HC value. In Appendix C, for a more detailed analysis of the distribution of probability masses between OOD and ID texts, we consider the logarithms of the probabilities of chosen and rejected sequences, similar to the toy example from Appendix B.
@@
As highlighted by Wang et al. (2023), models with higher alignment tend to produce less diverse generations. The dependency of HC metrics on Self-BLEU (Zhu et al., 2018) is shown in Appendix Figure 18 and is similar to the dependency on KL divergence. Based on these graphs, we affirm that the TR methods show higher HC metric values at the same level of response diversity (or KL divergence).
@@
Thus, simply lowering the $\beta$ coefficient is insufficient to achieve comparable results. By updating the reference policy, TR methods achieve higher metric values and maintain stable performance compared to standard alignment techniques, indicating the importance of incorporating such update strategies to enhance training dynamics. TR methods, therefore, achieve better HC values, although at higher KL divergences from the SFT policy. These results support our initial motivation for introducing the method.